OpenStudy (anonymous):

Determine the value(s) of the variable that must be excluded from the domain. x2^ - 16/-9x2^ + 27x + 36

OpenStudy (anonymous):

@dpaInc (:

OpenStudy (anonymous):

@abb50 Can you pleeaseee help me? (:

OpenStudy (anonymous):

@KNicksfan4life Omg, help me.

OpenStudy (anonymous):

\[x ^{2}-16/-9x ^{2}+27x+36\]

OpenStudy (anonymous):

x² - 16 / -9x² + 27x + 36 = (x-4)(x+4) / -9 (x² - 3x - 4) = (x-4)(x+4) / -9 (x-4)(x+ 1) = (x+4) / -9(x+ 1) ; x ≠ -1, 4 D: {xeR | x ≠ -1 , 4} ----------> the restriction are - 1 and 4